# EDBench: Large-Scale Electron Density Data for Molecular Modeling

### v1

@@ A.4 Significance of benchmark tasks
A.5 Detailed statistics of 6 benchmarks
We provide a detailed statistical analysis of six benchmarks in the EDBench suite: ED5-EC, ED5-OE, ED5-MM, ED5-OCS, ED5-MER, and ED5-EDP. Figures S2, S3, and S4 illustrate the distributions of the number of atoms, the number of ED points at the threshold $\rho_{\tau}=0$, and the per-molecule mean ED values at $\rho_{\tau}=0$, respectively. As shown, the number of ED points significantly exceeds the number of atoms, which provides richer information for force field learning and related downstream tasks.
@@ A.5 Detailed statistics of 6 benchmarks
A.6 Details of computational efficiency
@@
Next, we report the GPU memory usage and training time for each model. Due to varying memory requirements across models, we had to use different GPU devices to accommodate specific models and avoid out-of-memory (OOM) issues. Tables S2 and S3 present the computational efficiency of PointVector and X-3D, respectively. As expected, both GPU memory consumption and training time increase consistently with the number of sampling points $\xi$.
@@
A.7 Details of retrieval tasks (ED5-MER)
@@
[…]
@@
A.8 Details about the quality analysis of ED outputs generated by generation task
@@
A.9 Details of ablation study
@@
Ablation Study on the Number of Point Cloud Samples $\xi$. Point cloud-based methods (e.g., PointVector and X-3D) commonly adopt farthest point sampling (FPS) (Yu et al., 2022) to reduce the number of input points. Therefore, the number of sampled points, denoted as $\xi$, is a critical hyperparameter that directly affects both the model’s capacity to capture spatial structures and its computational efficiency. A larger number of points allows the model to better represent the geometric details of ED, particularly in regions with ambiguous boundaries or sharp density gradients, facilitating the learning of fine-grained spatial features. However, increasing $\xi$ also leads to higher memory consumption and longer training and inference times, especially when dealing with large-scale ED datasets. Therefore, choosing an appropriate number of points is essential to balance representational power and computational cost. To investigate this trade-off, we evaluate the performance of the point cloud-based PointVector model under different sample sizes $\xi=\{512,1024,2048,4096,8192\}$. Table S8 reports the results on the ED5-OE dataset. We observe that PointVector achieves the best performance when $\xi=2048$, reaching an average MAE of 0.0248. Additionally, the model performance does not monotonically improve with increasing $\xi$. This may be attributed to the model’s limited capacity—PointVector contains only 1.5454M parameters—which may constrain its ability to effectively leverage a large number of ED points. This observation highlights the need for more strong and ED-specialized architectures in future work.
Table S8: Ablation study (ED5-OE dataset) of PointVector on the number of sampling points $\xi$ with MAE×$\times 100$ metric and $\rho_{\tau}=0.05$.

### v2

@@ A.11 Significance of benchmark tasks
A.12 Detailed statistics of 6 benchmarks
We provide a detailed statistical analysis of six benchmarks in the EDBench suite: ED5-EC, ED5-OE, ED5-MM, ED5-OCS, ED5-MER, and ED5-EDP. Figures S2, S3, and S4 illustrate the distributions of the number of atoms, the number of ED points at the threshold $\rho_{\tau}=0$, and the per-molecule mean ED values at $\rho_{\tau}=0$, respectively. Specifically, ED length refers to the total number of ED sampling points retained after applying a density threshold $\rho_{\tau}$ (with $\rho_{\tau}=0$ meaning all ED values are retained). This metric is analogous to the commonly used notion of molecular length (i.e., the number of heavy atoms), and serves to quantify the effective representational size of the ED modality. As shown, the number of ED points significantly exceeds the number of atoms, which provides richer information for force field learning and related downstream tasks.
@@ A.12 Detailed statistics of 6 benchmarks
A.13 Details of computational efficiency
@@
Next, we report the GPU memory usage and training time for each model. Due to varying memory requirements across models, we had to use different GPU devices to accommodate specific models and avoid out-of-memory (OOM) issues. Tables S6 and S7 present the computational efficiency of PointVector and X-3D, respectively. As expected, both GPU memory consumption and training time increase consistently with the number of sampling points $\xi$.
@@
A.14 Training with full corpus
Firstly, we clarify the usage of the EDBench corpus in our benchmark construction. In the ED5-MER task, 50,000 molecules serve as anchors. For each anchor, we sample 10 negative molecules, resulting in a total of  550,000 molecules involved in ED5-MER alone. After deduplication across all benchmark tasks, the entire benchmark suite involves approximately 680,000 unique molecules, accounting for roughly 20% of the full corpus.
Next, we have conducted additional large-scale training experiments. Specifically, we replace each task’s training set with a pre-training corpus comprising about 2.67 million molecules not used in the benchmark. We use the pre-training corpus to train the model with an estimated training time of about two days, then evaluate it on the original validation and test sets of each task. Table S9 shows the performance of X-3D on ED5-OE. We observe that after training the model on the remaining large-scale dataset (X-3D (full)), its performance on the test set has been further improved. The results highlight the value of the full corpus for representation learning.
@@ A.14 Training with full corpus
MAE HOMO-2 HOMO-1 HOMO-0 LUMO+0 LUMO+1 LUMO+2 LUMO+3
X-3D (full) 1.5797 1.6359 1.9104 2.9981 2.7028 2.8725 2.8708
A.15 Effectiveness of extension to periodic systems
We conducted preliminary explorations to assess whether the data representations and model architectures proposed in EDBench can be effectively extended to periodic systems, such as crystalline solids.
Dataset. we extracted approximately 2,600 material molecules from the Materials Project (which primarily contains periodic systems, especially crystalline solids) and used the more suitable SCAN functional and Def2-SVP basis set to compute the electron density.
Metrics. We then split the dataset into training, validation, and test sets in an 8:1:1 ratio and evaluated it with tasks analogous to those in EDBench, including orbital energies prediction (EDMaterial-OE), multipole moment prediction (EDMaterial-MM), and electron density prediction from molecular structures (EDMaterial-EDP).
Experimental results. The experimental results are shown in Table S10 and Table S11. We find that X-3D and PointVector achieve strong performance on EDMaterial-OE and EDMaterial-MM, indicating that the proposed data representations and model architecture perform well in understanding material-based electron densities. This demonstrates that EDBench can be effectively extended to periodic systems, such as crystalline solids.
Table S10: MAE $\times 100$ performance of X-3D and PointVector on EDMaterial-OE task
Model HOMO-2 HOMO-1 HOMO-0 LUMO+0 LUMO+1 LUMO+2 LUMO+3
X-3D 3.45±0.02 3.47±0.01 3.21±0.02 3.41±0.02 4.04±0.02 4.39±0.04 4.53±0.07
PointVector 5.57±1.29 4.65±0.82 3.75±0.37 4.31±0.79 6.08±1.58 7.17±2.65 8.64±3.50
Table S11: MAE performance of X-3D and PointVector on EDMaterial-$M$ task.
Model Dipole X Dipole Y Dipole Z Magnitude
X-3D 0.8365±0.0098 1.0587±0.0280 1.2712±0.0577 1.3314±0.0187
PointVector 1.0568±0.0737 1.3476±0.0958 1.3615±0.0302 1.9749±0.0390
In addition, we find that DeepDFT is a structure-based method for generating electron densities and conduct further experiments using DeepDFT on both the ED5-EDP task and the newly constructed EDMaterial-EDP task. The results, presented in Table S12, show that DeepDFT achieved low MAE and high Pearson correlation for both ED5-EDP and EDMaterial-EDP. Notably, DeepDFT displayed lower performance in Spearman correlation, highlighting an area for future improvement.
Overall, these results further demonstrate the broader applicability of EDBench, especially in materials science, and we will continue to explore methods to improve performance, particularly in capturing non-linear relationships within electron-density fields.
Table S12: Performance of DeepDFT on EDMaterial-EDP task, showing MAE, Pearson, and Spearman correlations.
Model MAE Pearson Spearman
ED5-EDP 0.018±0.003 0.993±0.004 0.381±0.162
EDMaterial-EDP 0.118±0.029 0.918±0.034 0.633±0.115
A.16 Details of retrieval tasks (ED5-MER)
We first use GeoFormer and EquiformerV2 as molecular structure (MS) encoders, and PointVector and X-3D as electron density (ED) encoders. These encoders are combined pairwise—GeoFormer+PointVector, GeoFormer+X-3D, EquiFormer+PointVector, and EquiFormer+X-3D—to systematically evaluate cross-modal retrieval performance. Table S13 reports the Top-$k$ accuracy on both $\texttt{ED}\rightarrow\texttt{MS}$ and $\texttt{MS}\rightarrow\texttt{ED}$ tasks. Results reveal substantial performance differences among combinations. For example, GeoFormer+PointVector achieves only 17.67% Top-1 accuracy, while GeoFormer+X-3D reaches 68.32%, yielding an absolute improvement of 50.65%. Similarly, EquiFormer+PointVector achieves just 10.24% Top-1 accuracy, whereas EquiFormer+X-3D reaches 78.71%—an absolute gain of 68.47%. These results highlight the critical importance of selecting appropriate encoder architectures for effective cross-modal representation learning between MS and ED.
Table S13: The Top-$k$ accuracy (%) on ED5-MER dataset. $\texttt{ED}\rightarrow\texttt{MS}$ represents using electron density (ED) to retrieve molecular structure (MS).
MS model ED model $\texttt{ED}\rightarrow\texttt{MS}$ $\texttt{MS}\rightarrow\texttt{ED}$
@@ A.16 Details of retrieval tasks (ED5-MER)
A.17 Sensitivity analysis on the hardness of negative samples in retrieval tasks
Table S14: Retrieval Performance of ED5-MER with EquiformerV2 and X-3D Across Different Negative Sample Hardness (Random, Easy, Hard).
Split ED→MS MS→ED
Top-1 Top-2 Top-3 Top-1 Top-2 Top-3
Random 0.882 ± 0.005 0.959 ± 0.004 0.982 ± 0.001 0.882 ± 0.005 0.961 ± 0.001 0.983 ± 0.002
Easy 0.878 ± 0.004 0.957 ± 0.002 0.983 ± 0.002 0.873 ± 0.004 0.959 ± 0.001 0.983 ± 0.002
Hard 0.842 ± 0.009 0.947 ± 0.004 0.980 ± 0.001 0.836 ± 0.004 0.935 ± 0.005 0.971 ± 0.003
To explore how sensitive the ED5-MER performance is to the hardness of negative samples, we conducted retrieval experiments using a combination of the EquiformerV2 and X-3D models with negative samples of varying difficulty. Specifically, for each anchor, we used three types of negative samples, with five negative samples per type, meaning we needed to correctly retrieve the anchor from a set of six samples. The three types were as follows:
  • Random: Negative samples were randomly selected from the remaining dataset.
  • Easy: Negative samples were selected from a different cluster based on ECFP4 fingerprints and 3D USR descriptors (as detailed in the manuscript).
  • Hard: Negative samples were selected from the same cluster as the anchor.
The experimental results are summarized in Table S14. We observed that the "random" setup yielded the highest retrieval performance, followed by "easy" and "hard" in that order. This suggests that the difficulty of the retrieval task increases from random to easy to hard, indicating that ED5-MER performance is sensitive to the hardness of negative samples.
A.18 Error statistics for element-wise and size-resolved resolution
Here, we present element-wise and size-resolved error statistics, showing the MSE metrics for different electron density (ED) thresholds (0.1, 0.15, 0.2) across various atomic length ranges and atom types in the ED5-EDP task. As shown in Tables S15 and S16, we observe performance differences between these categories, which provide valuable insights into areas where the models may struggle. These results offer guidance for future research aimed at improving model performance in specific contexts.
Table S15: MSE metrics for different ED thresholds $\rho_{\tau}$ across atomic length ranges in the ED5-EDP task. MAX represents the maximum atomic length.
#atoms $\rho_{\tau}=0.1$ $\rho_{\tau}=0.15$ $\rho_{\tau}=0.2$
(0, 10) 1.4533 0.1127 0.0617
(10, 20) 2.2329 0.1581 0.0657
(20, 30) 1.8257 0.0615 0.0394
(30, 40) 2.0581 0.271 0.1009
(40, MAX) 1.2474 0.0093 0.0251
Table S16: MSE metrics for different ED thresholds $\rho_{\tau}$ across atom types in the ED5-EDP task.
atom type $\rho_{\tau}=0.1$ $\rho_{\tau}=0.15$ $\rho_{\tau}=0.2$
H 1.3387 0.0113 0.0262
B 0.6879 0.2677 2.069
C 1.0201 0.0063 0.0185
N 1.7475 0.0156 0.0234
O 3.0661 0.0551 0.0479
F 7.6878 0.1462 0.1472
Si 32.557 1.4758 0.7707
P 7.3819 2.8219 0.7615
S 27.2228 9.9254 3.5864
Cl 6.3195 0.7083 0.2215
Ge 85.3257 61.3484 3.5796
Se 3.3337 2.0152 0.5184
Br 85.2498 9.8684 8.6048
A.19 Details about the quality analysis of ED outputs generated by generation task
@@
A.20 Details of ablation study on threshold $\rho$ and the number of sampling $\xi$
@@
Ablation Study on the Number of Point Cloud Samples $\xi$. Point cloud-based methods (e.g., PointVector and X-3D) commonly adopt farthest point sampling (FPS) [83] to reduce the number of input points. Therefore, the number of sampled points, denoted as $\xi$, is a critical hyperparameter that directly affects both the model’s capacity to capture spatial structures and its computational efficiency. A larger number of points allows the model to better represent the geometric details of ED, particularly in regions with ambiguous boundaries or sharp density gradients, facilitating the learning of fine-grained spatial features. However, increasing $\xi$ also leads to higher memory consumption and longer training and inference times, especially when dealing with large-scale ED datasets. Therefore, choosing an appropriate number of points is essential to balance representational power and computational cost. To investigate this trade-off, we evaluate the performance of the point cloud-based PointVector model under different sample sizes $\xi=\{512,1024,2048,4096,8192\}$. Table S19 reports the results on the ED5-OE dataset. We observe that PointVector achieves the best performance when $\xi=2048$, reaching an average MAE of 0.0248. Additionally, the model performance does not monotonically improve with increasing $\xi$. This may be attributed to the model’s limited capacity—PointVector contains only 1.5454M parameters—which may constrain its ability to effectively leverage a large number of ED points. This observation highlights the need for more strong and ED-specialized architectures in future work.
[…]
@@
A.21 Details of ablation study on Bohr grid spacing
The choice of a fixed 0.4 Bohr grid spacing and density thresholding could potentially affect the representation of diffuse electron density, particularly around anionic species or Rydberg states. Therefore, we conducted an ablation study to evaluate the sensitivity of our models to different grid spacing settings, thereby strengthening the analysis in Section 4.5. Specifically, we generated additional versions of the ED5-OE dataset using three alternative grid spacings: 0.2, 0.3, and 0.5 Bohr. We then re-evaluated two representative models—PointVector and X-3D—on each of these datasets. The results, summarized in Table S20, show that PointVector achieves MAEs ranging from 0.0248 to 0.0277 (a maximum difference of 0.0029), and X-3D achieves MAEs between 0.0245 and 0.0259 (a maximum difference of 0.0014). These consistently narrow ranges indicate that both models are highly robust to the choice of grid spacing, with only marginal variations in MAE across the different resolutions.
This insensitivity can be attributed to two main factors: (i) Electron density changes smoothly in space, so even when a slightly coarser grid is used, the key local patterns in the density are still well preserved. As a result, grid-based representations can capture the essential features without significant loss of information; (ii) The downstream architectures (PointVector and X-3D) are designed to learn from spatial context and may be less reliant on fine-grained grid resolution once essential geometric and density features are retained. Moreover, while extremely diffuse states (e.g., Rydberg orbitals) may suffer from truncation in coarse grids, the EDBench primarily consists of compact, drug-like molecules where electron density is well localized. This further mitigates the impact of grid coarseness in the current setting.
[…]
Model Bohr HOMO-2 HOMO-1 HOMO-0 LUMO+0 LUMO+1 LUMO+2 LUMO+3 Mean
PointVector 0.2 1.94±0.11 1.77±0.04 2.02±0.04 3.51±0.21 3.39±0.26 3.47±0.19 3.31±0.16 2.77
0.3 1.80±0.03 1.74±0.03 1.97±0.03 3.23±0.11 3.05±0.13 3.24±0.04 3.14±0.02 2.60
0.4 1.73±0.01 1.68±0.01 1.92±0.01 3.08±0.05 2.86±0.05 3.05±0.02 3.01±0.02 2.48
0.5 1.84±0.05 1.78±0.07 2.00±0.06 3.37±0.23 3.17±0.25 3.35±0.22 3.23±0.15 2.68
X-3D 0.2 1.76±0.02 1.71±0.02 1.93±0.01 3.09±0.02 2.88±0.02 3.05±0.01 3.00±0.01 2.49
0.3 1.73±0.02 1.68±0.01 1.92±0.01 3.04±0.01 2.81±0.01 3.00±0.02 3.00±0.03 2.45
0.4 1.75±0.02 1.72±0.02 1.98±0.00 3.21±0.01 3.02±0.02 3.25±0.04 3.20±0.03 2.59
0.5 1.76±0.01 1.74±0.01 1.99±0.01 3.21±0.01 3.01±0.01 3.21±0.01 3.16±0.00 2.58
A.22 Details of ablation study on temperature $\tau$
Table S21: Retrieval performance of ED5-MER with EquiformerV2 and X-3D across different temperature $\tau=\{0.05,0.07,0.1,0.25,0.3,0.5\}$.
[…]
Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
0.05 0.796±0.006 0.952±0.004 0.982±0.002 0.790±0.003 0.947±0.003 0.980±0.002
0.07 0.787±0.007 0.948±0.004 0.981±0.001 0.784±0.007 0.942±0.001 0.977±0.003
0.1 0.782±0.009 0.943±0.005 0.977±0.002 0.782±0.006 0.939±0.004 0.976±0.001
0.25 0.776±0.002 0.939±0.001 0.975±0.001 0.775±0.003 0.937±0.002 0.972±0.001
0.3 0.776±0.007 0.938±0.001 0.974±0.002 0.778±0.004 0.935±0.003 0.971±0.001
0.5 0.775±0.002 0.940±0.001 0.976±0.000 0.777±0.002 0.934±0.001 0.970±0.001
In retrieval tasks, we used the fixed temperature $\tau$ = 0.07 in the InfoNCE loss, we chose this value based on empirical evidence from prior works. Specifically, $\tau$ = 0.07 has been shown to yield strong performance in contrastive learning frameworks [84, 64], and remains widely adopted in subsequent models [85, 86]. To assess robustness across temperature configurations, we conducted an ablation study varying $\tau\in\{0.05,0.1,0.25,0.3,0.5\}$, guided by values explored in prior literature [87, 88, 89, 90]. The experimental results are shown in Table S21. Overall, the retrieval performance across different temperatures is relatively consistent, with a maximum performance difference of just 2.1%. Additionally, we observed a clear trend: retrieval performance improves as the temperature decreases. For instance, in the ED→MS task, as the temperature decreases from 0.5 to 0.05, the Top-1 performance steadily increases from 0.775 to 0.796. Therefore, we recommend using a lower temperature to maintain higher performance.